Jacob wants to have $10,000 in his savings account in three years.  If he currently does not have anything in his account, how much will he need to save each month if his account compounds at 4% annually? please solve using excel
Answer #1

We can use excel function PMT to find the montly payment

PMT parameters: rate, nper, PV, FV

rate = Monthly effective rate = (1 + 0.04)^(1/12) - 1 = 0.003273739782

nper = 3 * 12 = 36

PV = 0

FV = 10000

Use excel function,PMT

PMT = -$262.18443925

(Ignore the negative sign)

Monthly payment = $$262.18443925

Can you please upvote? Thank You :-)